# Dr.Mi-Bench: A Modular-integrated Benchmark for Scientific Deep Research Agent

### v1

@@ 3.4 Efficiency Evaluation
Beyond correctness, a crucial aspect of any practical agent is its operational efficiency. Our evaluation protocol assesses this from two complementary perspectives: the trade-off between performance and processing time, and the relationship between the length of the final report and its generation time. First, to analyze the equivalent performance-time efficiency, we measure the end-to-end processing time (latency) in seconds for each task. This relationship is analyzed using Pareto Frontier analysis, which allows us to identify agents that offer an optimal balance between speed and accuracy and citation, distinguishing them from solutions that are merely fast but inaccurate, or accurate but slow. Second, we evaluate the generative efficiency of the agents: for each generated report $R$, we measure its total length in tokens. By correlating the token count with the processing time, we assess the agent’s generation throughput (tokens per second). This metric provides practical insights into the verbosity, computational cost, and user perceived latency of the different systems.
4 Dr.Mi-Bench Annotation
Our annotation methodology is fundamentally grounded in top-notch academic papers, which provide the implicit yet structured clues necessary to annotate objective evaluation metrics for core competencies of an agent: Planning, Retrieval, and Reasoning. Based on this principle, a team of trained domain experts (senior Ph.D. candidates) manually annotate a dataset of 200 instances across 10 diverse scientific disciplines. The overall framework is illustrated in Figure 1, while the appendices provide detailed annotation guidelines (Appendix A.2), quality management protocols (Appendix A.3), and complete examples for both a research paper and a review paper (Appendix A.4).
@@ 4.1 Data Sources
We ground our benchmark in peer-reviewed scientific papers, as their authenticity and quality provide an ideal source of test data for evaluating DR agents designed to emulate human experts.
@@
The curated papers were selected based on three primary criteria to ensure quality, diversity, and fairness. (i) Recency and Impact: All papers, including any pre-prints, were required to be publicly available after 2024 and have at least 10 citations. (ii) Task Diversity: The dataset contains a balanced distribution of both research papers, which test an agent’s ability to conduct a deep analysis of a specific research question, and review papers, which assess its ability to perform a broad synthesis of an entire research area. (iii) Disciplinary Fairness: The dataset is equally distributed across 10 disciplines, including materials, finance, chemistry, computer science, medicine, biology, environmental science, energy, building and construction, and Earth science. (iv) Non-STEM and Applied Scope: Beyond traditional core sciences, the benchmark explicitly incorporates non-STEM and applied domains, specifically Finance, Medicine, and Building and Construction, to ensure broad generalization capabilities. (v) Cross-Disciplinary Nature: We prioritized domains that inherently bridge multiple specialties (e.g., Materials and Finance) to assess the agents’ capacity for interdisciplinary knowledge synthesis.
@@
Our semi-automated curation process involved first using the Web of Science database to filter a candidate pool by discipline, date, and citations. Domain experts then manually selected the final papers from this pool, prioritizing quality, broad sub-field coverage, and the inclusion of landmark research from top-tier venues like Nature and ICLR.
@@ 4.2 Query Annotation
The first step in our annotation process is the creation of a realistic research query $Q$ for each source paper, formulated based on its abstract. Each query is designed to capture the paper’s core scientific contribution while strictly omitting paper-specific identifiers (e.g., author names or novel method acronyms), ensuring it represents an open-ended question a practitioner might realistically ask.
@@ 4.4 Annotation of Retrieval Module
The annotation of the gold-standard evidence set, $\mathcal{E}^{+\ast}$, is designed to be objective and tailored to the task’s goal of evaluating either depth or breadth. Accordingly, for research tasks, the source paper itself is designated as the sole piece of gold evidence. For review tasks, the paper’s complete bibliography constitutes the ground-truth evidence set. This list of evidence, comprising identifiers such as DOIs, arXiv IDs, book titles, and webpage links, was compiled via a semi-automated process using the CrossRef and Semantic Scholar APIs, supplemented with manual curation for any missing entries to ensure comprehensive and verifiable coverage.
@@ 5 End-to-End Experiments
Table 2: End-to-end performance across modules on research, review, and average tasks. All results are presented as percentages.
Research Review Avg. Model Accuracy Precision Recall F1 Accuracy Precision Recall F1 Accuracy Precision Recall F1 Planning Gemini 23.50 23.50 23.02 22.95 27.64 27.64 29.01 27.72 25.57 25.57 26.02 25.33 Retrieval OpenAI 62.00 0.69 1.73 5.27 1.36 31.35 Gemini 72.00 1.55 4.00 7.64 3.00 36.77 Perleixity 66.00 3.64 22.85 9.01 6.88 34.82 Grok 76.00 3.62 13.22 37.51 6.84 39.81 Reasoning OpenAI 60.72 61.58 40.96 48.27 56.58 61.59 38.71 46.63 58.65 61.59 39.83 47.45 Gemini 61.31 60.54 41.20 48.05 58.21 60.81 39.75 47.27 59.76 60.67 40.48 47.66 Perleixity 44.57 60.51 29.80 38.96 44.39 61.27 30.22 39.30 44.48 60.89 30.01 39.13 Grok 31.93 54.03 21.48 29.76 30.90 58.07 21.56 30.44 31.42 56.05 21.52 30.10 Search-r1 25.49 35.26 22.69 27.61 23.02 27.53 16.38 20.54 24.26 31.40 19.53 24.08
We assess state-of-the-art, end-to-end DR systems: OpenAI’s o3-deep-research (openai-blog), Google’s Gemini-2.5-pro Deep Research (gemini-dr), Perplexity Deep Research (perplexity-dr), Grok Deeper Search (grok-dr), and Search-r1 (jin2025searchr1trainingllmsreason). We employ GPT-4o (openai2023gpt4o) to automate the comparison between model outputs and our ground-truth data. To mitigate potential evaluation bias, we also performed a consistency check using a diverse set of LLMs (Appendix A.5).
@@ 5.1 What Are the Modular Strengths and Weaknesses of End-to-End DR Agents?
The main results, presented in Table 2 and Table 11, reveal that end-to-end scientific DR remains a significant challenge for state-of-the-art agent. The different DR agents exhibit specialized strengths, with different agents excelling in different core modules, while overall performance indicates considerable opportunity for future advancements. We provide a qualitative analysis of high- and low-scoring cases in Appendix A.9 to offer deeper insights into specific model behaviors and failure modes.
@@
The task of breaking down a research question into sub-tasks proves to be extremely difficult. Among the evaluated agents, Gemini is the only one with an explicit planning module that exposes its sub-tasks for evaluation. It achieves a modest total F1 score of only 25.33, highlighting the ability of scientific research question decomposition.
@@
In the Reasoning module, models demonstrate relatively stronger capabilities, though scores are still moderate. A clear top tier emerges, with Gemini and OpenAI delivering the best performance. A nuanced trade-off is visible between them: Gemini consistently leads in overall Accuracy (59.76) and Recall (40.48), while OpenAI excels in Precision (61.59). In stark contrast, models like Grok and Perplexity lag significantly in this area, indicating that strong retrieval skills do not necessarily translate to effective reasoning and synthesis.
@@ 5.2 How Does Agent Performance Differ Across Specific Domains?
Our analysis of end-to-end DR agents reveals a clear performance hierarchy dependent on the scientific domain (Figure 2). We consistently observe the highest performance in fields like Medicine and Biology, which is likely attributable to the extensive representation of these domains in the models’ training corpora. In contrast, agents struggle significantly more in specialized areas such as Finance and Materials, which are likely less prevalent in the training data. This performance gap indicates that the combined challenge of retrieval and reasoning varies considerably across different disciplines. This performance disparity reveals a significant bias towards data-rich fields and calls for a concerted research effort to bridge the capability gap for these underperforming but equally important scientific disciplines.
@@
5.3 What Are the Performance-Cost Trade-offs in Deep Research Agents?
We analyze the efficiency of each end-to-end agent by examining its two core modules: retrieval and reasoning, aiming to maximize accuracy while minimizing latency and output tokens. This analysis, framed by the concept of the Pareto Frontier, reveals completely different efficiency landscapes within these critical stages of the DR process, as shown in Figure 3.
@@ 5.3 What Are the Performance-Cost Trade-offs in Deep Research Agents?
The retrieval module’s efficiency landscape does not present a balanced trade-off. Grok is the sole agent on the Pareto frontier, as it dominates other systems by being both significantly faster (~76s) and more accurate (39.81%) in retrieval, rendering competitors Pareto-inefficient (Figure 3(a)). In contrast, the reasoning module exhibits a classic speed-accuracy trade-off, with all evaluated agents lying on the Pareto Frontier (Figure 3(b)). This creates a balanced curve where users can choose between the highest-accuracy model, OpenAI (~59% Acc), and the fastest option, Search-r1 (~30s).
@@
Our analysis of the relationship between reasoning accuracy and report length (token count) reveals a complex, non-linear trade-off. A strong positive correlation between verbosity and accuracy holds for most models, but the top performers show a key difference in token efficiency. As shown in Figure 3(c), Gemini delivers the highest accuracy (~59.76%) from a concise ~13k token report, while OpenAI is nearly twice as verbose for no performance gain. This suggests that Gemini is significantly more token-efficient and that excessive verbosity does not guarantee superior performance. Below this top tier, other models like Perplexity and Grok exhibit a simpler positive correlation where shorter reports are progressively less accurate.
@@
5.4 Does Our Metric Effectively Align with Human Preferences?
To validate our automatic metrics against human judgment, we conducted a human evaluation study where domain experts performed pairwise comparisons on 100 blinded report pairs. Experts selected the superior report based on an overall assessment of its factuality, comprehensiveness, and coherence. Our primary finding is a high degree of alignment: the ranking produced by our automatic Reasoning F1 score matched the human expert’s preference in 86.23% of the comparisons. This strong agreement validates our benchmark as a reliable proxy for what experts consider a high-quality research report.
Model Accuracy Precision Recall F1
Retrieval
Post-Cutoff 25.27
Full 25.18
Reasoning
Post-Cutoff 59.41 60.18 39.34 46.66
Full 59.94 59.30 39.74 46.56
Model Accuracy Precision Recall F1
Retrieval
Post-Cutoff 37.15
Full 36.77
Reasoning
Post-Cutoff 63.97 60.97 43.33 49.91
Full 59.76 60.67 40.48 47.66
5.5 Do Agents Overestimate Their Capabilities through Data Leakage?
Forward-citation Leakage.
"Forward citations" (papers published after the source) could distort evaluation metrics by artificially lowering retrieval scores or inflating reasoning capabilities through information leakage. To address this, we implement a temporal filtering experiment by explicitly adding publication date constraints to the agent prompts. As shown in Table 4 (detailed in Table 9), the overall performance shift is minimal (Retrieval Accuracy from 25.18% to 25.27%), confirming that forward citation leakage has a negligible impact and validating the robustness of our results.
Cut-off Dates Leakage.
5.6 Is the Benchmark Saturated to Test-Time Compute?
@@
The benchmark is effectively unsaturated. We conducted pass@k experiments (k=1, 3) on a 20-instance subset. As shown in Figure 4 (detailed in Table 10), a consistent performance gain from pass@1 to pass@3 across all modules. It indicates that the benchmark successfully measures the benefit of increased test-time compute and is not merely limited by rigid evaluation formats. Even with increased compute (pass@3), the highest average Reasoning score is 68.49% (OpenAI), and Planning remains around 40%. This substantial gap from 100% confirms that the benchmark is effectively unsaturated and remains a challenging standard for future, more capable agents.
@@ 6 Isolated Experiments
This section extends our analysis to the foundational LLMs, assessing their potential as the core backbones for DR agents. This includes prominent models such as GPT-o3 (gpt-o3), Gemini-2.5-Pro (gemini-2.5-family), Perplexity-sonar-pro (perplexity-sonar-pro), Grok-4 (grok-4), Claude-3.7-sonnet (claude-3-7-sonnet), as well as models from the Llama (grattafiori2024llama) (Llama-3.2-1B/3B-Instruct, Llama-3.1-8B-Instruct) and Qwen Instruct (team5qwen2) (Qwen2.5-1.5B/3B/7B-Instruct) series. To isolate their intrinsic capabilities, our analysis bypasses the retrieval module, often a separate, tool-based component.
@@ 6.1 Is High-Level Planning a Key Bottleneck for Foundational LLMs?
Research Review Avg. Model Accuracy Precision Recall F1 Accuracy Precision Recall F1 Accuracy Precision Recall F1 GPT-o3 27.24 27.24 24.10 25.52 29.48 29.48 25.70 27.37 28.36 28.36 24.90 26.45 Gemini-2.5-pro 31.31 31.31 26.14 28.42 31.22 31.22 25.91 28.25 31.27 31.27 26.02 28.34 Sonar-reasoning-pro 29.41 29.41 26.13 27.58 26.93 26.93 23.56 25.08 28.17 28.17 24.85 26.33 Grok-4 23.75 23.75 20.73 22.08 29.98 29.98 26.59 28.10 26.87 26.87 23.66 25.09 Claude-3.7-sonnet 24.12 24.12 21.83 22.86 26.15 26.15 24.60 25.29 25.14 25.14 23.21 24.07
@@ 6.2 What is the Performance Hierarchy Among LLMs for Reasoning Module?
Research Review Avg. Model Accuracy Precision Recall F1 Accuracy Precision Recall F1 Accuracy Precision Recall F1 GPT-o3 85.14 63.70 56.91 59.70 59.72 60.11 40.79 47.78 72.43 61.91 48.85 53.74 Gemini-2.5-pro 69.09 61.69 46.88 52.79 87.17 63.02 57.90 59.99 78.13 62.36 52.39 56.39 Sonar-reasoning-pro 83.91 62.98 55.44 58.33 62.55 62.13 42.45 49.71 73.23 62.56 48.95 54.02 Grok-4 82.89 63.50 55.04 58.38 54.43 60.88 37.16 45.17 68.66 62.19 46.10 51.77 Claude-3.7-sonnet 84.76 63.17 56.52 59.22 61.56 61.66 41.86 49.15 73.16 62.42 49.19 54.18
Our analysis of the reasoning module reveals a clear performance hierarchy, where Gemini-2.5-pro distinguishes itself from a pack of otherwise closely-performing models through its unique ability to handle complex, broad-scope documents. As shown in Table 6, Gemini-2.5-pro leads with an average F1 score of 56.39. The key capability that separates the top performer from the rest is its performance on complex review papers. Gemini-2.5-pro is the only model whose F1 score significantly improves on these broader tasks (rising from 52.79 on research to 59.99 on review). In stark contrast, every other model in the evaluation experiences a notable performance degradation on review papers, with F1 scores dropping by 8-13 points. We provide a detailed comparison of open- and closed-source LLMs in Appendix A.7. This drop is primarily driven by a sharp fall in Recall, suggesting that while most models can reason effectively over a focused source, they struggle to synthesize the wide-ranging information presented in a literature review.

### v2

@@ 3.4 Efficiency Evaluation
Beyond correctness, a crucial aspect of any practical agent is its operational efficiency. Our evaluation protocol assesses this from two complementary perspectives: the trade-off between performance and processing time, and the relationship between the length of the final report and its generation time. First, to analyze the equivalent performance-time efficiency, we measure the end-to-end processing time (latency) in seconds for each task. This relationship is analyzed using Pareto Frontier analysis on end-to-end latency to identify the optimal trade-off between speed, token, price and performance. Second, we evaluate the generative efficiency of the agents: for each generated report $R$, we measure its total length in tokens. By correlating the token count with the processing time, we assess the agent’s generation throughput (tokens per second). This metric provides practical insights into the verbosity, computational cost, and user perceived latency of the different systems.
4 ADRA-Bank Annotation
Our annotation methodology is fundamentally grounded in top-notch academic papers, which provide the implicit yet structured clues necessary to annotate objective evaluation metrics for core competencies of an agent: Planning, Retrieval, and Reasoning. Based on this principle, a team of trained domain experts (senior Ph.D. candidates) manually annotate a dataset of 200 instances across 10 diverse academic disciplines. The overall framework is illustrated in Figure 1, while the appendices provide detailed annotation statistics (Appendix H.1), guidelines (Appendix H.2), quality management protocols (Appendix J), and complete examples for both research and review papers (Appendix H.3).
@@ 4.1 Data Sources
We ground our benchmark in peer-reviewed academic papers, leveraging their authenticity and quality to evaluate DR agents designed to emulate human experts.
@@
We selected papers based on three criteria to ensure quality, diversity, and fairness. (i) Recency and Impact: All papers, including any pre-prints, were required to be publicly available after 2024 and $\geq 10$ citations. (ii) Task Diversity: The dataset balances research papers, which conduct deep analysis of a specific research question, and review papers, which perform broad synthesis of an entire research area. (iii) Disciplinary Fairness: Instances are equally distributed across 10 disciplines: materials, finance, chemistry, computer science, medicine, biology, environmental science, energy, building and construction, and earth science. Appendix D details selection bias and cross-disciplinary factors.
@@
Our semi-automated curation process involved first using the Web of Science database to filter a candidate pool by discipline, date, and citations. Domain experts then manually selected the final papers from this pool, prioritizing quality, broad sub-field coverage, and the inclusion of landmark research from top-tier venues.
@@ 4.2 Query Annotation
The first step in our annotation process is the creation of a realistic research query $Q$ for each source paper, formulated based on its abstract. Each query is designed to capture the paper’s core academic contribution while strictly omitting paper-specific identifiers (e.g., author names or novel method acronyms), ensuring it represents an open-ended question a practitioner might realistically ask.
@@ 4.4 Annotation of Retrieval Module
The gold-standard evidence set $\mathcal{E}^{+\ast}$is tailored to evaluate either depth or breadth. Accordingly, for research tasks, the source paper itself is designated as the sole piece of gold evidence. For review tasks, the paper’s complete bibliography constitutes the ground-truth evidence set. This list of evidence, comprising identifiers such as DOIs, arXiv IDs, book titles, and webpage links, was compiled via a semi-automated process using the CrossRef and Semantic Scholar APIs, supplemented with manual curation for any missing entries to ensure comprehensive and verifiable coverage.
@@ 5 End-to-End Experiments
We assess state-of-the-art, end-to-end DR systems: OpenAI’s o3-deep-research (OpenAI, 2025), Google’s Gemini-2.5-pro Deep Research (Google, ), Perplexity Deep Research (PerplexityAI, ), Grok Deeper Search (xAI, ), and Search-r1 (Jin et al., 2025b). We employ GPT-4o (OpenAI, 2024) to automate the comparison between model outputs and our ground-truth data. To mitigate potential evaluation bias, we also performed a consistency check using a diverse set of LLMs (Appendix K.4).
@@ 5.1 What Are the Modular Strengths and Weaknesses of End-to-End DR Agents?
The main results, presented in Table 2, reveal that end-to-end academic DR remains a significant challenge for state-of-the-art agent. DR agents excel in different core modules, yet overall performance leaves significant room for improvement. We provide a qualitative analysis of high- and low-scoring cases in Appendix K.8 to offer deeper insights into specific model behaviors and failure modes.
@@
The task of breaking down a research question into sub-tasks proves to be extremely difficult. Among the evaluated agents, Gemini is the only one with an explicit planning module that exposes its sub-tasks for evaluation. It achieves a modest total F1 score of only 25.33, highlighting the ability of academic research question decomposition.
@@
In the Reasoning module, models show stronger yet moderate capabilities. A clear top tier emerges, with Gemini and OpenAI delivering the best performance. A nuanced trade-off is visible between them: Gemini consistently leads in overall Accuracy (59.76) and Recall (40.48), while OpenAI excels in Precision (61.59). In stark contrast, models like Grok and Perplexity lag significantly in this area, indicating that strong retrieval skills do not necessarily translate to effective reasoning and synthesis.
@@ 5.2 How Does Agent Performance Differ Across Specific Domains?
Our analysis of end-to-end DR agents reveals a clear performance hierarchy dependent on the academic domain (Figure 2). We observe the highest performance in fields like Medicine and Biology, which is likely attributable to the extensive representation of these domains in the models’ training corpora. In contrast, agents struggle significantly more in specialized areas such as Finance and Materials, which are likely less prevalent in the training data. This performance gap indicates that the combined challenge of retrieval and reasoning varies considerably across different disciplines. This performance disparity reveals a significant bias towards data-rich fields and calls for a concerted research effort to bridge the capability gap for these underperforming but equally important academic disciplines.
@@
5.3 What Are the Performance-Cost Trade-offs in DR?
We analyze the efficiency of end-to-end agent by examining two core modules: retrieval and reasoning, aiming to maximize accuracy while minimizing latency, output tokens and cost (Appendix K.3). The Pareto Frontier analysis reveals completely different efficiency landscapes within these critical stages of the DR process, as shown in Figure 3.
@@ 5.3 What Are the Performance-Cost Trade-offs in DR?
The retrieval module’s efficiency landscape does not present a balanced trade-off. Grok is the sole agent on the Pareto frontier, as it dominates other systems by being both significantly faster (~76s) and more accurate (39.81%) in retrieval, rendering competitors Pareto-inefficient (Figure 3(a)). In contrast, the reasoning module exhibits a classic speed-accuracy trade-off, with all evaluated agents lying on the Pareto Frontier (Figure 3(b)). This creates a balanced curve where users can choose between the highest-accuracy model, OpenAI (~59% Accuracy), and the fastest option, Search-r1 (~30s).
@@
Our analysis of the relationship between reasoning accuracy and report length reveals a non-linear trade-off. A strong positive correlation between verbosity and accuracy holds for most models, but the top performers show a key difference in token efficiency. As shown in Figure 3(c), Gemini delivers the highest accuracy (~59.76%) from a concise ~13k token report, while OpenAI is nearly twice as verbose for no performance gain. This suggests that Gemini is significantly more token-efficient and that excessive verbosity does not guarantee superior performance. Below this top tier, other models like Perplexity and Grok exhibit a simpler positive correlation where shorter reports are progressively less accurate.
@@
5.4 Does Metric Align with Human Preferences?
To validate our metrics against human judgment, domain experts performed pairwise comparisons on 100 blinded report pairs, assessing factuality, comprehensiveness, and coherence. Experts selected the superior report based on an overall assessment of its factuality, comprehensiveness, and coherence. We find a high degree of alignment: our Reasoning F1 score matched expert preferences in 86.23% of cases. This strong agreement confirms the benchmark as a reliable proxy for expert-judged research quality.
Table 3: End-to-end performance with gold plans across modules on research, review, and average. Values are reported as percentages.
[…]
5.5 Is Planning as the Dominant Bottleneck?
To identify deep research pipeline bottlenecks, we conducted a Gold Plan Injection ablation study, providing agents with ground-truth plans while keeping model-generated retrieval and reasoning. Results in Table 3 show that fixing the planning module significantly boosts Reasoning accuracy. Specifically, OpenAI’s accuracy rose from 58.65% (end-to-end) to 70.93%, a 12% gain confirming planning as a dominant bottleneck; high-quality plans effectively guide reasoning even with imperfect retrieval.
5.6 Is ADRA-Bank Saturated to Test-Time Compute?
@@
The benchmark is effectively unsaturated. Pass@k experiments ($k=1,3$) on a 20-instance subset (Figure 4, Table 7) show consistent gains across all modules from pass@1 to pass@3. This confirms the benchmark effectively measures test-time compute benefits rather than being limited by rigid formats. Even at pass@3, OpenAI’s top Reasoning score is only 68.49%, while Planning hovers around 40%. This substantial headroom proves the benchmark remains a rigorous and unsaturated challenge for future agents.
[…]
@@ 6 Isolated Experiments
This section extends our analysis to the foundational LLMs, assessing their potential as the core backbones for DR agents. This includes prominent models such as GPT-o3 (OpenAI, ), Gemini-2.5-Pro (Google, ), Perplexity-sonar-pro (PerplexityAI, ), Grok-4 (OpenAI, ), Claude-3.7-sonnet (Claude, ), as well as models from the Llama (Grattafiori et al., 2024) (Llama-3.2-1B/3B-Instruct, Llama-3.1-8B-Instruct) and Qwen Instruct (Team, 2024) (Qwen2.5-1.5B/3B/7B-Instruct) series. To isolate their intrinsic capabilities, our analysis bypasses the retrieval module, often a separate, tool-based component.
@@ 6.2 What is the Performance Hierarchy Among LLMs for Reasoning Module?
Our analysis of the reasoning module reveals a clear performance hierarchy, where Gemini-2.5-pro distinguishes itself from a pack of otherwise closely-performing models through its unique ability to handle complex, broad-scope documents. As shown in Table 5, Gemini-2.5-pro leads with an average F1 score of 56.39. The key capability that separates the top performer from the rest is its performance on complex review papers. Gemini-2.5-pro is the only model whose F1 score significantly improves on these broader tasks (rising from 52.79 on research to 59.99 on review). In stark contrast, every other model in the evaluation experiences a notable performance degradation on review papers, with F1 scores dropping by 8-13 points. We provide a detailed comparison of open- and closed-source LLMs in Appendix K.6. This drop is primarily driven by a sharp fall in Recall, suggesting that while most models can reason effectively over a focused source, they struggle to synthesize the wide-ranging information presented in a literature review.